# Graph-Based Bidirectional Transformer Decision Threshold Adjustment Algorithm for Class-Imbalanced Molecular Data

### v1

@@ 4 Conclusion
Given the prevalence of imbalanced molecular data sets, where the class sizes vary vastly, the ability of an algorithm to accurately predict the class of data elements in imbalanced data sets is extremely important. However, many classification algorithms do not perform well on highly imbalanced data as they tend to overpredict the majority class. This paper presents a BTDT-MBO method for molecular data classification in case of imbalanced data sets with highly varied class sizes. The algorithm, incorporating Merriman-Bence-Osher (MBO) techniques as well as a bidirectional transformer, integrates adjustments in the classification threshold for machine learning classifiers in order to deal with the class imbalance. To demonstrate the utility of the proposed method, particular attention is given to highly imbalanced data sets. Numerical experiments show that the proposed BTDT-MBO method performs very well even with high class imbalance ratios; in particular, the computational experiments demonstrate that the results of the proposed method are almost always superior to those of the comparison techniques. Overall, the new model serves as a powerful machine learning tool for data sets with a high class imbalance.
The present work additionally investigates distance correlation as a weight function in the proposed BTDT-MBO method, serving as an alternative to the Gaussian kernel used to compute graph weights. The results for the BTDT-MBO algorithm using a distance correlation weight function are similar to those using Gaussian weights, with the distance correlation model performing better for one data set; these results suggest further utility of the distance correlation in imbalanced molecular data classification and similarity graph-based methods. In particular, data that is far from a normal distribution may benefit from the use of a distance correlation weight function rather than a Gaussian weight function.
Future work includes exploring imposing class size constraints into the technique. Moreover, we would like to explore integrating other correlation techniques, as well as incorporating extended-connectivity fingerprints or autoencoders into the algorithm structure.
Acknowledgements
This work is supported in part by NSF grant DMS-2052983. The work of Guo-Wei Wei was supported in part by NIH grants R01GM126189, R01AI164266 and R35GM148196, NSF grants DMS-2052983, DMS-1761320, and IIS-1900473, NASA grant 80NSSC21M0023, Michigan Economic Development Corporation, MSU Foundation, Bristol-Myers Squibb 65109, and Pfizer.

### v3

@@ 4 Conclusion
Given the prevalence of molecular data where the class sizes vary vastly, the ability of an algorithm to accurately predict the class of data elements in imbalanced data sets is extremely important. However, many classification techniques are not accurate on highly imbalanced data as they tend to overestimate majority classes. This paper presents a BTDT-MBO method for molecular data classification in case of imbalanced data sets with highly varied class sizes. The method, which integrates Merriman-Bence-Osher procedures as well as a bidirectional transformer, incorporates adjustments in the data classification threshold for machine learning classifiers to handle the high class imbalance. To illustrate the advantages of the proposed technique, particular attention is given to highly imbalanced data sets.
Numerical experiments indicate that the BTDT-MBO procedure performs very well even with high class imbalance ratios; in particular, the computational experiments on DS1 data sets and DrugMatrix data sets demonstrate that the results of the proposed procedure are almost always more accurate than the comparison algorithms such as 1, as indicated by Figures 2 and 3. The ROC-AUC score was used as a metric. Overall, the new model serves as a powerful machine learning tool for data sets with a high class imbalance.
The present work additionally investigates the distance correlation as a choice for a weight function in the proposed BTDT-MBO algorithm, serving as an alternative to the Gaussian kernel often used to compute graph weights. Overall, the results for the BTDT-MBO technique using a distance correlation weight function are similar to those using Gaussian weights, with the distance correlation model performing better for one data set; these results suggest further utility of the distance correlation in imbalanced molecular data classification and similarity graph-based methods. In particular, data that is far from a normal distribution may benefit from the use of a distance correlation weight function rather than a Gaussian weight function.
Funding Information
The work of this paper is supported partly by the NSF DMS-2052983 grant. The third author was supported partly by NIH R01GM126189, R01AI164266 and R35GM148196 grants, NSF grants DMS-1761320, DMS-2052983, and IIS-1900473, NASA grant 80NSSC21M0023, Michigan Economic Development Corporation, MSU Foundation, Pfizer, and Bristol-Myers Squibb 65109.
Conflict of Interest
All authors declare that they have no conflicts of interest.
Author Contributions
Nicole Hayes was responsible for the software, computational experiments and writing. Ekaterina Rapinchuk and Guo-Wei Wei were responsible for methodology, reviewing, editing, some writing and supervision.
Data Availability
The data sets in this paper are publically available at https://github.com/rinikerlab/GHOST/tree/main/data.
Statement of Usage of Artificial Intelligence
The authors have not received any help from artificial intelligence such as ChatGPT while preparing this paper.